Close to the center of a campus is a tall silo topped with a hemispherical cap. The cap is
frictionless when wet. Someone has somehow balanced a pumpkin at the highest point. The
line from the center of curvature of the cap to the pumpkin makes an angle
1
Expert's answer
2020-06-17T09:24:40-0400



Let initially pumpkin was at top then it slides to point P.


Applying energy conservation law,

Energy at top = Energy at point P.

Energy at top is only due to it's potential energy.

Energy at point P is due to it's kinetic energy and potential energy.


So,

"mgR = mgRcos\theta + \frac {mv^2}{2}"


"2gR(1-cos\theta) = v^2" . . . . . . . . . . . (i)



Now, at point P, balancing forces so that it keep contact with hemisphere.


"mgcos\theta = \frac {mv^2}{R}"


"gRcos\theta = v^2" . . . . . . . . . . . . (ii)



solving (i) and (ii)


"2gR(1-cos\theta) = gRcos\theta"


"2 - 2cos\theta = cos\theta"


"2 = 3cos\theta\implies cos\theta = (\frac{2}{3})"


"\theta" is angle with vertical.


Numerically, "\theta = 48.2^0"